Answer the following:

a.     MPC = .7.What is the government spending multiplier?

= 1/1-0.7 = 10/3 = 3.33

b. MPC = .85.What is the tax multiplier?

= -(0.85/1-0.85) = -17/3 = -5.67

c. If the government spending multiplier is 5, what is the tax multiplier?

d. If the tax multiplier is -3, what is the government spending multiplier?

e. If government purchases and taxes are increased by $150 billion simultaneously, what will the effect be on equilibrium output (income)?
Answer #1

1) Government spending multiplier is given as

1/1-MPC=1/1-0.7=1/0.3=3.34

2)Tax multiplier= -MPC/MPS=-0.85/1-0.85=-0.85/0.15=(-)5.67

3) Government multiplier=1/1-MPC

5=1/1-MPC

5-5MPC=1

5Mpc=4

MPC=4/5=0.8

MPS=1-0.8=0.2

So tax multiplier= -MPC/MPS= -0.8/0.2=4

d) Tax multiplier=-MPC/MPS=-3

-MPC/MPS=-3/1

MPC/MPS=3/1

So in th ratio form MPC is 3 times MPS. Or in other words MPC=0.75 and MPS is 0.25

Government spending multiplier=1/1-mpc=1/1-0.75=1/0.25=4

e) IF both government purchase and taxes are simultaneously increased, then income because of balanced budget would also increase by $150 billion. Output would increase by $150 billion